OpenStudy (anonymous):

how do i multiply 16 radical 8 times radical 216

OpenStudy (anonymous):

\[16\sqrt{8} \times \sqrt{216}\]

OpenStudy (amistre64):

216=6.6.6 for starters

OpenStudy (amistre64):

8 = 2.2.2

OpenStudy (amistre64):

sqrt(6.2.6.2.6.2) = 6.2 = 12

OpenStudy (amistre64):

12sqrt(12) that is ...forgot to finish it lol

OpenStudy (amistre64):

\[16\sqrt{8}\sqrt{216}=16\sqrt{12^3}=16*12\sqrt{12}=16*12\sqrt{4*3}=16*12*2\sqrt{3}\] maybe

OpenStudy (amistre64):

\[a^nb^n=(ab)^n\] \[\sqrt{a}\sqrt{b}=a^{1/2}b^{1/2}=(ab)^{1/2}=\sqrt{ab}\]

OpenStudy (anonymous):

i can't see the end of that second to last post, can you write it again?